# LLM Security Guard for Code

### v1

@@ Abstract.
Many developers rely on Large Language Models (LLMs) to facilitate software development. Nevertheless, these models have exhibited limited capabilities in the security domain. We introduce LLMSecGuard, an open-source framework that offers enhanced code security through the synergy between static code analyzers and LLMs. LLMSecGuard aims to equip practitioners with code solutions that are more secure than the code initially generated by LLMs. It also benchmarks LLMs, providing valuable insights into the evolving security properties of these models.
@@
journalyear: 2024conference: 28th International Conference on Evaluation and Assessment in Software Engineering; June 18–21, 2024; Salerno, Italybooktitle: 28th International Conference on Evaluation and Assessment in Software Engineering (EASE 2024), June 18–21, 2024, Salerno, Italycopyright: nonedoi: 10.1145/xxxccs: Security and privacy Software security engineering
@@ 1. Introduction
We introduce LLMSecGuard, a framework designed to extensively examine the security properties of LLMs and leverage their analytical capabilities for secure code development. It applies static security analysis on LLM-generated code to uncover potential security issues and guides LLMs in resolving such issues in the code. Moreover, LLMSecGuard assesses the security properties of LLMs and benchmark them across different CWEs. In summary, LLMSecGuard contributes to more secure software development. Unlike existing code assistants, developers can integrate unlimited LLMs and code analysis engines into this framework through REST APIs. LLMSecGuard is open-source and publicly available on GitHub. [^1] https://github.com/aryakvnust/LLMSecGuard
@@ 21 String combined = salt + username + password;
ChatGPT identified the constant key, and it correctly suggested the secure “GCM” encryption mode as a best practice. However, violations such as “static salt for key generation” and “hardcoded password” were in the response without any proper warning. The suggested code, at line 23, included the “SHA-256” hash function for key generation, which is not a best practice. Anyway, when we asked about the security risk at line 6 (i.e., static salt), ChatGPT responded: “Using a hard-coded string as a salt is generally not considered a good practice for security. Ideally, salt should be unique for each user and should be randomly generated for each password. Using a constant salt for all users could make the system more vulnerable to certain types of attacks, such as rainbow table attacks”. It then provided the code example in Listing 2 as a replacement.
@@ 3. LLMSecGuard
We introduce LLMSecGuard, an open-source framework that offers enhanced code security through the synergy between code security analyzers and LLMs. The primary objective of LLMSecGuard is (i) to enhance the security of LLM-generated code, and (ii) to benchmark the security properties of LLMs. It adopts a RESTful architecture, implemented in Python using Django and Flask, and uses SQLite for data persistence. LLMSecGuard is publicly available on GitHub. [^3] https://github.com/aryakvnust/LLMSecGuard
@@
Figure 1 illustrates the schematic architecture of LLMSecGuard, comprising three main components: “Prompt Agent”, “Security Agent”, and “Benchmark Agent”.
@@
Benchmark Agent. This component puts different LLMs to security test. Particularly, it evaluates the security properties of LLMs based on a set of challenges, structured in JSON format. Each challenge includes a prompt and the expected outcome, and “Benchmark Agent” is responsible for flagging LLMs that pass the challenge.
@@ 4.3. Secure Code Generation
Figure 2 illustrates the workflow in this scenario. It begins with the user providing a prompt of interest, anticipating the output to be a piece of code. The “Prompt Agent” assumes control and forwards the user’s prompt to the top LLM determined in the benchmarking scenario (unless the user chooses a different LLM). Upon receiving the LLM’s response, this component then transfers the produced code to the “Security Agent” for security inspection. The “Security Agent” collaborates with external analysis engines (configured earlier) to uncover potential vulnerabilities and respective lines in the code. If the termination condition is met (i.e., there is no vulnerability or the maximum analysis iterations is reached), the code along with its vulnerability information, is immediately reported back to the user. In cases where the termination condition is not met, this information is relayed to the “Prompt Agent”. This component formulates a new prompt based on the collected vulnerability information and queries the LLM once again.
@@ 7. Conclusion
We introduced LLMSecGuard, an open-source framework developed to equip developers with code solutions that are more secure than the code initially generated by Large Language Models (LLMs). These code suggestions are obtained through the integration of LLMs and static security code analyzers. LLMSecGuard also measures the security properties of LLMs and provides researchers and practitioners with a current security comparison of different LLMs in the wild.

### v2

# Preprint Version

@@ Abstract.
Many developers rely on Large Language Models (LLMs) to facilitate software development. Nevertheless, these models have exhibited limited capabilities in the security domain. We introduce LLMSecGuard, a framework to offer enhanced code security through the synergy between static code analyzers and LLMs. LLMSecGuard is open source and aims to equip developers with code solutions that are more secure than the code initially generated by LLMs. This framework also has a benchmarking feature, aimed at providing insights into the evolving security attributes of these models.
@@
journalyear: 2024copyright: rightsretainedconference: 28th International Conference on Evaluation and Assessment in Software Engineering; June 18–21, 2024; Salerno, Italybooktitle: 28th International Conference on Evaluation and Assessment in Software Engineering (EASE 2024), June 18–21, 2024, Salerno, Italydoi: 10.1145/3661167.3661263isbn: 979-8-4007-1701-7/24/06copyright: nonedoi: 10.1145/xxxccs: Security and privacy Software security engineering
@@ 1. Introduction
We introduce LLMSecGuard, a framework specifically designed to examine the security properties of LLMs for code generation and harness their analytical capabilities to facilitate secure code development. It applies static security analysis on LLM-generated code to uncover potential security issues and guides LLMs in resolving such issues in the code. Moreover, LLMSecGuard can assess the security properties of LLMs and benchmark them across different CWEs. In summary, LLMSecGuard contributes to more secure software development. Unlike existing code assistants, developers can integrate unlimited LLMs and code analysis engines into this framework through REST APIs. LLMSecGuard is open-source and publicly available on GitHub. [^1] https://github.com/aryakvnust/LLMSecGuard
@@ 21 String combined = salt + username + password;
ChatGPT identified the constant key, and it correctly suggested the secure “GCM” encryption mode as a best practice. However, violations such as “static salt for key generation” and “hardcoded password” were in the response without any proper warning. The suggested code, at line 23, included the “SHA-256” hash function for key generation, which is not a best practice. Anyway, when we asked about the security risk at line 6 (ie., static salt), ChatGPT responded: “Using a hard-coded string as a salt is generally not considered a good practice for security. Ideally, salt should be unique for each user and should be randomly generated for each password. Using a constant salt for all users could make the system more vulnerable to certain types of attacks, such as rainbow table attacks”. It then provided the code example in Listing 2 as a replacement.
@@ 3. LLMSecGuard
We introduce LLMSecGuard, an open-source framework to offer enhanced code security through the synergy between code security analyzers and LLMs. The primary objective of LLMSecGuard is (i) to enhance the security of LLM-generated code, and (ii) to benchmark the security properties of LLMs. It adopts a RESTful architecture, implemented in Python using Django and Flask, and uses SQLite for data persistence. Figure 1 illustrates the schematic architecture of LLMSecGuard. We explain the main components in the following.
@@
Benchmark Agent. This component puts different LLMs to security test. Particularly, it evaluates the security properties of LLMs based on a set of challenges structured in JSON format. Each challenge includes a prompt and the expected outcome, and “Benchmark Agent” is responsible for flagging LLMs that pass the challenge.
@@ 4.3. Secure Code Generation
Figure 2 illustrates the workflow in this scenario. It begins with the user providing a prompt of interest, anticipating the output to be a piece of code. The “Prompt Agent” assumes control and forwards the user’s prompt to the top LLM determined in the benchmarking scenario (unless the user chooses a different LLM). Upon receiving the LLM’s response, this component then transfers the produced code to the “Security Agent” for security inspection. The “Security Agent” collaborates with external analysis engines (configured earlier) to uncover potential vulnerabilities and respective lines in the code. If the termination condition is met (i.e., there is no vulnerability or the maximum analysis iterations is reached), the code, along with its vulnerability information, is immediately reported back to the user. In cases where the termination condition is not met, this information is relayed to the “Prompt Agent”. This component formulates a new prompt based on the collected vulnerability information and queries the LLM once again.
@@ 7. Conclusion
We introduced LLMSecGuard, an open-source framework developed to equip developers with code solutions that are more secure than the code initially generated by Large Language Models (LLMs). These code suggestions are obtained through the integration of LLMs and static security code analyzers. LLMSecGuard also offers a feature to measure the security properties of LLMs and provide a current security comparison of different LLMs in the wild.